Displaying page layout variables

Kent McPherson a.o.^{∗}

1995/11/23

1 Introduction

This L^{A}TEX 2ε package is a reimplementation of layout.sty by Kent McPherson. It defines the command \layout which produces an overview of the layout of the current document. The command \layout* recomputes the values it uses to produce the overview.

∗Converted for L^{A}TEX 2ε by Johannes Braams